# JHU Applied Physics Laboratory, 11100 Johns Hopkins Rd, Laurel, MD 20723, USA

### v1

@@ 4.1 Light curve fitting with PACMAN
We ran a Monte-Carlo-Marcov-Chain (MCMC) with the Python package emcee (Foreman-Mackey et al., 2013) to obtain the final parameter estimates and uncertainties. We used broad, uninformed linear priors for all parameters of the systematics model and the transit depth. We ran chains with 100000 steps and 500 walkers, removing the first 4000 steps of each chain as burn-in. This ensured that the chains were longer than 50 times the autocorrelation time for astrophysical and systemic parameters and that the chains converged. The uncertainties on the parameters were derived from the posterior distributions and based on the Gaussian deviations of the parameters from their median value. Minor correlations are seen in the posterior distributions of the constant flux offsets and the linear slopes of flux with time. Furthermore, the row shift slopes between forward and reverse scans are strongly correlated. Since both parameters describe the shift of the spectral trace on the detector, a tight correlation between these two systemic corrections is expected. The obtained transit depth is not strongly correlated with the other parameters.
@@
From the broadband light curve, we derive a transit depth of HD 86226 c of ppm in the full 1.1–1.7 band observed with HST/WFC3. This results in a planet radius of , assuming the stellar radius given in Table 5. This planet radius is slightly larger than the value of derived by Teske et al. (2020). While the planet may appear smaller in the wavelength range covered by TESS, the radius difference of less than doesn’t allow for definite conclusions. For the analysis of the planet spectrum in Sect. 3, we assumed a planet radius of , as derived from the HST data.
@@ 4.2 Light curve fitting with Eureka!
In this equation, are free parameters, and is the time since the visit started. Eureka! furthermore provides the capability to fit a second-order polynomial per orbit with the parameters , which we did not use. An additional parameter, , was varied as a fixed parameter with possible values of either , , or . It fits the timing offset of the hstramp model.
A range of other correction techniques were tested, including individually- and simultaneously-fitted transits, positional adjustments (with the inbuilt functions xpos, ypos, ywidth), and separate consideration of forward and reverse scan directions, where spectra generated by only single scan directions were evaluated. While effective under more stable conditions, these methods proved challenging to implement for this dataset. Isolating data segments by removing integrations or entire transits did not fully resolve the issues.
@@
Parameter estimates and uncertainties for the fits were obtained using a Monte-Carlo-Marcov-Chain (MCMC) of light curve fits with the python package emcee (Foreman-Mackey et al., 2013). The final light curve fits still show a large contribution of red noise in most of the 27 spectral bins between and . A major contributor to this systematic noise is the inclusion of the first orbit of every visit. The exponential ramp of this orbit is not identical to the ramp in the subsequent orbits, even after removing different amounts of exposures at the start of this orbit. Unfortunately, removing the first orbit of each visit in this reduction was impossible, as the many half-covered transits terminated the fitting with Eureka! in this case. The transit depth of the first wavelength bin between and deviates strongly from the remaining bins, which is why we remove it from further analysis.
@@ 4.3 Differences between the spectra obtained with PACMAN and Eureka!
Another difference between the systematics models is how we account for the upstream-downstream effect and the associated flux dependency on the spectral trace position. While PACMAN jointly fits all data with a constant offset between forward and reverse scans of a visit, Eureka! fits the scan directions separately. Including more data in a single fit might be beneficial for robustly determining the parameters of the complex underlying model. The row shift correction implemented in PACMAN significantly improved the light curve fits and reduced the scatter in the derived spectrum. However, the derived spectroscopic transit depths deviate from the uncorrected values by less than . This correction therefore only partially accounts for the differences in the spectra.
@@
These differences in the reduction also lead to different qualities of the light curve fits: while the RMS of the PACMAN spectroscopic light curve fits typically is only about 10% above the photon noise, the Eureka! RMS is about twice the photon noise in many bins. Based on these findings, we consider the PACMAN results more reliable than those from Eureka!. PACMAN was designed specifically for HST/WFC3, and this is a particularly challenging data set. The main problems in the Eureka! analysis arose from the bright host star and corresponding poor trace position stability and from the partial transit observations. While these effects were accounted for with PACMAN, Eureka! does not currently provide the required capabilities. Given the featureless nature of the spectrum, implementing these capabilities to Eureka! is out of the scope of this analysis. For further analysis and modeling of the planetary atmosphere, we will proceed with the spectrum obtained with the PACMAN pipeline.
@@ 5 Transit spectrum models
The featureless spectrum strongly rules out a clear solar-metallicity atmosphere for HD 86226 c, but other scenarios such as high atmospheric metallicity or high altitude clouds are still possible (Benneke & Seager, 2013). To quantify these potential scenarios, we used the retrieval package petitRADTRANS (pRT, Mollière et al., 2019; Nasedkin et al., 2024) to fit atmospheric models to the spectrum of HD 86226 c (see Sect. 5.1). In addition, we ran self-consistent cloud and haze models for the planet, which are described in Sections 5.2 and 5.3.
@@ 5.1 Retrievals with petitRADTRANS
The second approach assumed a simplified atmosphere consisting only of H2, He, and H2O. We focused on H2O because it has a strong feature in the observed wavelength range between and and is a species with high expected abundances based on planet formation models (Fortney et al., 2013; Bitsch et al., 2021; Izidoro et al., 2021; Burn et al., 2024). For a planet with such high equilibrium temperature (), we do not expect high abundances of CH4 as it will be photodissociated (Zahnle et al., 2009; Gao et al., 2020). Other species that could be expected in the atmosphere of HD 86226 c are CO and CO2 (e.g., Moses et al., 2013). However, these only have shallow features in the observed wavelength range, which is why we neglected them for this analysis. During the retrieval, we fitted for the H2O mass fraction, , and calculated the H and He mass fractions according to their solar ratio (see Asplund et al., 2009). We note that this ratio could differ for sub-Neptunes, as a preferential mass-loss of H over He could alter the atmospheric composition (Malsky et al., 2023). However, our data is not sensitive enough to determine a change in the H-He ratio.
@@ 5.2 Cloud forward models
We also explored cloud forward models of metal-enriched atmospheres to investigate plausible cloud compositions. To do this, we used the Thermal Stability Cloud (TSC) model (Blecic et al., submitted). This model was already successfully applied to multiple synthetic and actual HST and JWST targets to distinguish various cloud species present in their atmospheres and constrain the overall cloud structure (e.g., Venot et al., 2020; Taylor et al., 2023; Bell et al., 2024). TSC is a complex Mie-scattering cloud model that is built on the approaches of Benneke (2015) and Ackerman & Marley (2001), allowing for additional flexibility in the location of the cloud base (determined by the overall atmospheric metallicity) and in the width of the log-normal distribution (which affects the droplet sizes and distribution). The model formulation employs multiple free parameters, enabling us to capture the complex structure of the clouds. These include the cloud’s shape, location, and extent; the droplet sizes, distribution, and volume mixing ratio; and the overall atmospheric metallicity. This parametrized model assumes thermally stable equilibrium atmospheric conditions (Marley et al., 2013) and does not account for non-equilibrium microphysical processes. For more details on the model parameters, refer to the model description in Venot et al. (2020), as well as Ackerman & Marley (2001) and Benneke (2015).
@@
After obtaining the self-consistent temperature profiles for all atmospheric metallicities, we investigated which cloud species could condense within the HD 86226 c temperature regimes. These include the species KCl, ZnS, Na2S, MnS, SiO2, Cr, MgSiO3, Fe, and Mg2SiO4 (e.g., Gao et al., 2021; Morley et al., 2012a). We excluded KCl, ZnS, Na2S, Mg2SiO4, Cr, and SiO2 from our analysis for the following reasons: KCl, ZnS, and Na2S condense at lower temperatures and do not intersect the temperature-pressure () profiles for the metallicity range we considered, preventing cloud formation based on our thermal-stability approach. Mg2SiO4 condenses at slightly higher temperatures than MgSiO3. At the temperatures of HD 86226 c, MgSiO3 condensates will form first and be more abundant (Lodders & Fegley Jr, 2006). Cr is much less abundant than MnS or MgSiO3, significantly reducing the likelihood of Cr cloud formation (Lodders, 2003; Morley et al., 2012a). In the presence of both Mg and Si elemental species, Visscher et al. (2010) suggest that MgSiO3 forms before SiO2, making SiO2 less likely to form at all. Additionally, neither Cr nor SiO2 has prominent features at the wavelengths of our observations (Blecic et al., submitted), making them indistinguishable from other, more likely, cloud species.
@@
Fig. 8 shows the vapor-pressure () curves of the species we included in our analysis, which could condense between temperatures of 1000–3000 K and pressures of 100–10-8 bar. It is apparent that the position of the curves in the space strongly depends on the atmospheric metallicity (e.g., Morley et al., 2012a) shifting to lower temperatures for lower solar metallicity values and to higher temperatures for super-solar metallicities. We show this range as shaded regions for several cloud species. The plot demonstrates that the MnS, MgSiO3, and Fe curves intersect with the curves for our explored metallicity and pressure range. Even at a metallicity of 1000 solar, Na2S clouds could condense only at extremely low pressures, where their abundance is too low to produce visible signatures in the planetary spectra. Thus, we only proceeded with our grid exploration for the MnS, MgSiO3, and Fe clouds. In Fig. 9, we also present the scattering and absorption coefficients for these species. While MnS and MgSiO3 exhibit distinct features within the displayed wavelength range, they, along with Fe, are nearly indistinguishable in the range of our observations (1.1 – 1.7 m). This lack of differentiation has posed a challenge for our analysis, as discussed below.
@@
We generated a grid of cloud models across all metallicities, varying fundamental cloud parameters that can produce visible signatures in the spectra. In addition to metallicity, these include droplet radius, , and droplet volume mixing ratio, , described by the number of cloud droplets over the number of H2 molecules per unit volume. The presence of clouds effectively increases the radius of the planet and decreases the atmospheric scale height. A similar effect on the spectra could result from uncertainties in the planetary transit radius, potentially leading to either an underestimation or overestimation of the cloud droplet volume mixing ratio and droplet size. For this reason, we made the transit radius a free parameter in our model and we also allow for a vertical offset of the observed spectrum.
@@
With each parameter set, our cloud model calculated the cloud’s scattering and absorption cross-sections and optical depth. The refractive indices of MnS and MgSiO3 condensate species were taken from Wakeford & Sing (2015), while those for Fe are taken from Palik (1985). As a test, we also computed the MnS models with the refractive indices of Kitzmann & Heng (2018), which yielded qualitatively the same results. The scattering and absorption efficiencies were calculated using Mie-scattering theory (Toon & Ackerman, 1981) by utilizing the Mie-scattering code provided by Mark Marley (private communication). This code generates pre-calculated tables of scattering and absorption efficiencies over a range of wavelengths and droplet radii, on which we then interpolated depending on our wavelength region and the cloud droplet radius (see also Fig. 9).
@@
Fig. 10 shows some of the lowest reduced chi-squared () models and the grid exploration with the goodness of fit for MnS, MgSiO3, and Fe clouds. We explored the full range of our model parameters, but only the ranges that best fit the data are shown in the tables to reduce clutter. The tables list the minimum value for each set of parameters. The color bar range is set between 1 and 2 to emphasize only the models that best fit the data.
The top panels of Fig. 10 explore the MnS cloud models. In the top left panel, we show several clear, no-cloud models and some representative cloud models with the lowest . Only the flat model and high-metallicity models fit the data well (with reduced ), while lower metallicity models require some clouds to match the data. Note that the no-cloud model at 1solar metallicity requires a high planetary radius of = to match the data level. This value is much larger than the value our broadband light curve analysis reports, which confirms our initial assessment that this planet either has high-altitude clouds, which increase the planetary radius, or a high metallicity atmosphere, which effectively does the same.
The middle and right panels present the values for various models across the grid. To investigate whether there are trends when exploring the grid of our model parameters, these panels depict how changing each crucial cloud parameter (while keeping the others fixed at plausible values) influences the goodness of fit. Our analysis explored all combinations of cloud parameter values across our grid, but we chose to display panels with the most values smaller than two (shown as yellow to green colours in the panels). Other parameter combinations also occasionally yielded 2, but not as frequently as the selected examples.
@@
Fe grid model exploration leads to even more degeneracies. The spectra are flatter than for MnS clouds, but values are a bit larger than for MgSiO3 clouds. Fe models with values around one spread more uniformly across all metallicities, providing no clear indication of which atmospheric metallicities might be preferred, but suggesting that slightly larger droplet radius (on the order of 1-10m) and are needed to fit the data.
We also performed a full grid exploration assuming a planetary radius of 2.16 as given by Teske et al. (2020). These models, as models assuming other planetary radii, do not provide a good fit to the data. Teske et al. (2020)’s planetary radius is derived based on only optical observations and is smaller than our broadband light curve analysis radius of 2.313 . Models with the transit radius set to 2.16 have the base level of the spectrum at a much lower transit depth. To compensate for this and attempt to match the data (by lifting the spectrum and increasing the transit depth), the cloud droplet volume mixing ratio increases to high, less-plausible levels (on the order of to for super-solar metallicities above 30 solar), strongly biasing our results towards a low-metallicity solution.
Overall, the results show that we cannot conclusively identify which clouds condense in the atmosphere of HD 86226 c, or whether we have a clear, high metallicity atmosphere or an atmosphere with clouds. None of the considered species (MnS, MgSiO3, and Fe) exhibit distinct features within the wavelength range of our observations ( – ), which complicates the identification of unique cloud signatures in the data. If clouds are present, their effect would be more gray and uniform across all wavelengths, making it impossible to distinguish which particular species muted the spectral features of HD 86226 c. Just marginally, our cloudy solutions match the data better than the high metallicity, no-cloud solutions. The implications of these findings are further discussed in Section 6.

### v2

@@ 4.1 Light curve fitting with PACMAN
We ran a Marcov Chain Monte Carlo (MCMC) algorithm with the Python package emcee (Foreman-Mackey et al., 2013) to obtain the final parameter estimates and uncertainties. We used broad, uninformed linear priors for all parameters of the systematics model and the transit depth. We ran chains with 100000 steps and 500 walkers, removing the first 4000 steps of each chain as burn-in. This ensured that the chains were longer than 50 times the autocorrelation time for astrophysical and systemic parameters and that the chains converged. The uncertainties on the parameters were derived from the posterior distributions and based on the Gaussian deviations of the parameters from their median value. Minor correlations are seen in the posterior distributions of the constant flux offsets and the linear slopes of flux with time. Furthermore, the row shift slopes between forward and reverse scans are strongly correlated. Since both parameters describe the shift of the spectral trace on the detector, a tight correlation between these two systemic corrections is expected. The obtained transit depth is not strongly correlated with the other parameters.
@@
From the broadband light curve, we derive a transit depth of HD 86226 c of ppm in the full 1.1–1.7 band observed with HST/WFC3. This results in a planet radius of , assuming the stellar radius given in Table 5. This planet radius is slightly larger than the value of derived by Teske et al. (2020). While the planet may appear smaller in the wavelength range covered by TESS, the radius difference of less than does not allow for definite conclusions. For the analysis of the planet spectrum in Sect. 3, we assumed a planet radius of , as derived from the HST data.
@@ 4.2 Light curve fitting with Eureka!
In this equation, are free parameters, and is the time since the visit started. Eureka! furthermore provides the capability to fit a second-order polynomial per orbit with the parameters , which we did not use. An additional parameter, , was varied as a fixed parameter with possible values of , , or . It fits the timing offset of the hstramp model.
A range of other correction techniques were tested, including individually and simultaneously fitted transits, positional adjustments (with the inbuilt functions xpos, ypos, ywidth), and separate consideration of forward and reverse scan directions, where spectra generated by only single scan directions were evaluated. While effective under more stable conditions, these methods proved challenging to implement for this dataset. Isolating data segments by removing integrations or entire transits did not fully resolve the issues.
@@
Parameter estimates and uncertainties for the fits were obtained using a MCMC algorithm for the light curve fits with the python package emcee (Foreman-Mackey et al., 2013). The final light curve fits still show a large contribution of red noise in most of the 27 spectral bins between and . A major contributor to this systematic noise is the inclusion of the first orbit of every visit. The exponential ramp of this orbit is not identical to the ramp in the subsequent orbits, even after removing different amounts of exposures at the start of this orbit. Unfortunately, removing the first orbit of each visit in this reduction was impossible, as the many half-covered transits terminated the fitting with Eureka! in this case. The transit depth of the first wavelength bin between and deviates strongly from the remaining bins, which is why we remove it from further analysis.
@@ 4.3 Differences between the spectra obtained with PACMAN and Eureka!
Another difference between the systematics models is how we account for the upstream-downstream effect and the associated flux dependence on the spectral trace position. While PACMAN jointly fits all data with a constant offset between forward and reverse scans of a visit, Eureka! fits the scan directions separately. Including more data in a single fit might be beneficial for robustly determining the parameters of the complex underlying model. The row shift correction implemented in PACMAN significantly improved the light curve fits and reduced the scatter in the derived spectrum. However, the derived spectroscopic transit depths deviate from the uncorrected values by less than . This correction therefore only partially accounts for the differences in the spectra.
@@
These differences in the reduction also lead to different qualities of the light curve fits: While the RMS of the PACMAN spectroscopic light curve fits is typically only about 10% above the photon noise, the Eureka! RMS is about twice the photon noise in many bins. Based on these findings, we consider the PACMAN results more reliable than those from Eureka!. PACMAN was designed specifically for HST/WFC3, and this is a particularly challenging dataset. The main problems in the Eureka! analysis arose from the bright host star and corresponding poor trace position stability and from the partial transit observations. While these effects were accounted for with PACMAN, Eureka! does not currently provide the required capabilities. Given the featureless nature of the spectrum, implementing these capabilities to Eureka! is out of the scope of this analysis. For further analysis and modeling of the planetary atmosphere, we proceeded with the spectrum obtained with the PACMAN pipeline.
@@ 5 Transit spectrum models
The featureless spectrum strongly rules out a clear solar-metallicity atmosphere for HD 86226 c, but other scenarios such as high atmospheric metallicity or high altitude clouds are still possible (Benneke & Seager, 2013). To quantify these potential scenarios, we used the retrieval package petitRADTRANS (pRT; Mollière et al., 2019; Nasedkin et al., 2024) to fit atmospheric models to the spectrum of HD 86226 c (see Sect. 5.1). In addition, we ran self-consistent cloud and haze models for the planet, which are described in Sections 5.2 and 5.3.
@@ 5.1 Retrievals with petitRADTRANS
The second approach assumed a simplified atmosphere consisting only of H2, He, and H2O. We focused on H2O because it has a strong feature in the observed wavelength range between and and is a species with high expected abundances based on planet formation models (Fortney et al., 2013; Bitsch et al., 2021; Izidoro et al., 2021; Burn et al., 2024). For a planet with such a high equilibrium temperature (), we do not expect high abundances of CH4 as it is photodissociated (Zahnle et al., 2009; Gao et al., 2020). Other species that could be expected in the atmosphere of HD 86226 c are CO and CO2 (e.g., Moses et al., 2013). However, these only have shallow features in the observed wavelength range, which is why we neglected them for this analysis. During the retrieval, we fitted for the H2O mass fraction, , and calculated the H and He mass fractions according to their solar ratio (see Asplund et al., 2009). We note that this ratio could differ for sub-Neptunes, as a preferential mass-loss of H over He could alter the atmospheric composition (Malsky et al., 2023). However, our data is not sensitive enough to determine a change in the H-He ratio.
@@ 5.2 Cloud forward models
We also explored cloud forward models of metal-enriched atmospheres to investigate plausible cloud compositions. To do this, we used the Thermal Stability Cloud (TSC) model (Blecic et al., 2025). This model was already successfully applied to multiple synthetic and actual HST and JWST targets to distinguish various cloud species present in their atmospheres and constrain the overall cloud structure (e.g., Venot et al., 2020; Taylor et al., 2023; Bell et al., 2024). TSC is a complex Mie-scattering cloud model that is built on the approaches of Benneke (2015) and Ackerman & Marley (2001), allowing for additional flexibility in the location of the cloud base (determined by the overall atmospheric metallicity) and in the width of the log-normal distribution (which affects the droplet sizes and distribution). The model formulation employs multiple free parameters, enabling us to capture the complex structure of the clouds. These include the cloud’s shape, location, and extent; the droplet sizes, distribution, and volume mixing ratio; and the overall atmospheric metallicity. This parametrized model assumes thermally stable equilibrium atmospheric conditions (Marley et al., 2013) and does not account for non-equilibrium microphysical processes. For more details on the model parameters, refer to the model description in Venot et al. (2020), as well as Ackerman & Marley (2001) and Benneke (2015).
@@
After obtaining the self-consistent temperature profiles for all atmospheric metallicities, we investigated which cloud species could condense within the HD 86226 c temperature regimes. These include the species KCl, ZnS, Na2S, MnS, SiO2, Cr, MgSiO3, Fe, and Mg2SiO4 (e.g., Gao et al., 2021; Morley et al., 2012a). We excluded KCl, ZnS, Na2S, Mg2SiO4, Cr, and SiO2 from our analysis for the following reasons: KCl, ZnS, and Na2S condense at lower temperatures and do not intersect the temperature-pressure () profiles for the metallicity range we considered, preventing cloud formation based on our thermal-stability approach. Mg2SiO4 condenses at slightly higher temperatures than MgSiO3. At the temperatures of HD 86226 c, MgSiO3 condensates form first and be more abundant (Lodders & Fegley Jr, 2006). We note that Cr is much less abundant than MnS or MgSiO3, thus significantly reducing the likelihood of Cr cloud formation (Lodders, 2003; Morley et al., 2012a). In the presence of both Mg and Si elemental species, Visscher et al. (2010) suggest that MgSiO3 forms before SiO2, making SiO2 less likely to form at all. Additionally, neither Cr nor SiO2 has prominent features at the wavelengths of our observations (Blecic et al., 2025), making them indistinguishable from other, more likely, cloud species.
@@
Figure 8 shows the vapor-pressure () curves of the species we included in our analysis, which could condense between temperatures of 1000–3000 K and pressures of 100–10-8 bar. It is apparent that the position of the curves in the space strongly depends on the atmospheric metallicity (e.g., Morley et al., 2012a) shifting to lower temperatures for lower solar metallicity values and to higher temperatures for super-solar metallicities. We show this range as shaded regions for several cloud species. The plot demonstrates that the MnS, MgSiO3, and Fe curves intersect with the curves for our explored metallicity and pressure range. Even at a metallicity of 1000 solar, Na2S clouds could condense only at extremely low pressures, where their abundance is too low to produce visible signatures in the planetary spectra. Thus, we only proceeded with our grid exploration for the MnS, MgSiO3, and Fe clouds. In Fig. 9, we also present the scattering and absorption coefficients for these species. While MnS and MgSiO3 exhibit distinct features within the displayed wavelength range, they, along with Fe, are nearly indistinguishable in the range of our observations (1.1 – 1.7 m). This lack of differentiation has posed a challenge for our analysis, as discussed below.
@@
We generated a grid of cloud models across all metallicities, varying fundamental cloud parameters that can produce visible signatures in the spectra. In addition to metallicity, these include droplet radius, , and droplet volume mixing ratio, , described by the number of cloud droplets over the number of H2 molecules per unit volume. The presence of clouds effectively increases the radius of the planet and decreases the atmospheric scale height. A similar effect on the spectra could result from uncertainties in the planetary transit radius, potentially leading to either an underestimation or overestimation of the cloud droplet volume mixing ratio and droplet size. For this reason, we made the transit radius a free parameter in our model, and we also allowed for a vertical offset of the observed spectrum.
@@
With each parameter set, our cloud model calculated the cloud’s scattering and absorption cross-sections and optical depth. The refractive indices of MnS and MgSiO3 condensate species were taken from Wakeford & Sing (2015), while those for Fe were taken from Palik (1985). As a test, we also computed the MnS models with the refractive indices of Kitzmann & Heng (2018), which yielded qualitatively the same results. The scattering and absorption efficiencies were calculated using Mie-scattering theory (Toon & Ackerman, 1981) by utilizing the Mie-scattering code provided by Mark Marley (private communication). This code generates pre-calculated tables of scattering and absorption efficiencies over a range of wavelengths and droplet radii, on which we then interpolated depending on our wavelength region and the cloud droplet radius (see also Fig. 9).
@@
Figure 10 shows some of the lowest reduced chi-squared () models and the grid exploration with the goodness of fit for MnS, MgSiO3, and Fe clouds. We explored the full range of our model parameters, but only the ranges that best fit the data are shown in the tables to reduce clutter. The tables list the minimum value for each set of parameters. The color bar range is set between 1 and 2 to emphasize only the models that best fit the data.
The top panels of Fig. 10 explore the MnS cloud models. In the top left panel, we show several clear, no-cloud models and some representative cloud models with the lowest . Only the flat model and high-metallicity models fit the data well (with reduced ), while lower metallicity models require some clouds to match the data. We note that the no-cloud model at 1solar metallicity requires a high planetary radius of = to match the data level. This value is much larger than the value our broadband light curve analysis reports, which confirms our initial assessment that this planet either has high-altitude clouds, which increase the planetary radius, or a high metallicity atmosphere, which effectively does the same.
The middle and right panels present the values for various models across the grid. To investigate whether there are trends when exploring the grid of our model parameters, these panels depict how changing each crucial cloud parameter (while keeping the others fixed at plausible values) influences the goodness of fit. Our analysis explored all combinations of cloud parameter values across our grid, but we chose to display panels with the most values smaller than two (shown as yellow to green colors in the panels). Other parameter combinations also occasionally yielded 2, but not as frequently as the selected examples.
@@
The Fe grid model exploration led to even more degeneracies. The spectra are flatter than for MnS clouds, but values are a bit larger than for MgSiO3 clouds. The Fe models with values around one spread more uniformly across all metallicities, providing no clear indication of which atmospheric metallicities might be preferred but suggesting that a slightly larger droplet radius (on the order of 1-10m) and are needed to fit the data.
We also performed a full grid exploration assuming a planetary radius of 2.16 as given by Teske et al. (2020). These models, as models assuming other planetary radii, do not provide a good fit to the data. Teske et al. (2020)’s planetary radius is derived based on only optical observations and is smaller than our broadband light curve analysis radius of 2.313 . Models with the transit radius set to 2.16 have the base level of the spectrum at a much lower transit depth. To compensate for this and attempt to match the data (by lifting the spectrum and increasing the transit depth), the cloud droplet volume mixing ratio increases to high, less plausible levels (on the order of to for super-solar metallicities above 30 solar), strongly biasing our results toward a low-metallicity solution.
Overall, the results show that we cannot conclusively identify which clouds condense in the atmosphere of HD 86226 c, or whether we have a clear, high metallicity atmosphere or an atmosphere with clouds. None of the considered species (MnS, MgSiO3, and Fe) exhibit distinct features within the wavelength range of our observations ( – ), which complicates the identification of unique cloud signatures in the data. If clouds were present, their effect would be grayer and uniform across all wavelengths, which would make it impossible to identify the particular species that mutes the spectral features of HD 86226 c. Just marginally, our cloudy solutions match the data better than the high metallicity, no-cloud solutions. The implications of these findings are further discussed in Section 6.